OpenStudy (lazyboy):

give an example of a function. Give the relation as a set of ordered pairs. Use a test such as mapping or the verical line test to show if your set of ordered pairs is a function. An example of a function could be: distance traveled and time, or the cost of a turkey and its weight. You can list your relation as a set of ordered pairs, or as a table. In either case, you must prove the relation is a function

OpenStudy (lazyboy):

@nincompoop

OpenStudy (nincompoop):

try doing the function of time by distances traveled so in a graph: 

OpenStudy (lazyboy):

okay

OpenStudy (nincompoop):

functions are written in this format f(x)= some equation this is similar to y = mx + b the f(x) is the y read a little bit about functions http://finedrafts.com/files/math/precal/Larson%20PreCal%208th/Larson%20Precal%20CH2.pdf

OpenStudy (nincompoop):

so for example: a car is traveling 60 kilometers per hour (we need to label our graph accordingly). we need to be able to show how long will it take for a car to travel 35 miles, 45 miles and maybe 50 miles.

OpenStudy (lazyboy):

Okay, I see

OpenStudy (lazyboy):

Whats next?

OpenStudy (lazyboy):

well thanks for the help anyway! :)

OpenStudy (nincompoop):



OpenStudy (lazyboy):

Is that it or is that only a small step for the "big picture"

OpenStudy (nincompoop):

big picture. you need to be able to represent your equation graphically and mathematically

OpenStudy (nincompoop):

can you come up with a formula that determines the function of time?

OpenStudy (nincompoop):

we said that the the rate of speed is 60 km/hr so in minutes: that is 60 km/60 mins since 1 hour is 60 mins 60/60 is 1. So now, we have 1 km/min D=f(t)=1t if we want to use the rate rate per minute D=f(f)=60t if we want to use the rate per hour

OpenStudy (lazyboy):

I am sorry I was busy, now I am back

OpenStudy (nincompoop):

typo on my previous post it's D = f(t) = 60t

OpenStudy (lazyboy):

oh okay

OpenStudy (nincompoop):

so pretty much there's your equation and you can do your algebra to determine time or distance

OpenStudy (lazyboy):

okay, but I am not sure how I am supposed to submit my assignment. I mean like what information should I include and how should i write it

OpenStudy (nincompoop):

you can now apply it to the example we want to use

OpenStudy (nincompoop):

it's written just like how I've written them. you might want to rephrase. Did you read the link I sent you?

OpenStudy (lazyboy):

yes, I did

OpenStudy (lazyboy):

You wrote so much I am not sure what I am supposed to use

OpenStudy (nincompoop):

so what don't you understand?